RhondaSommer (rhondasommer):

Can someone show me how to do this? Write the equation of the sine function with amplitude of 7, period of 2π , phase shift of 90°, and a vertical shift of 15.

OpenStudy (plainntall):

What does all that mean? For the general formula for sine \[asin(bx-c) + d\] amplitude = |a| Period = 2pi/b Phase Shift = c/b Vertical shift = d so where do we go from there?

RhondaSommer (rhondasommer):

\_(o.o)_/ I have nooo idea

OpenStudy (sloppycanada):

Okay, so if you an amplitude of 7, that's more or less considered "a". asin(bx - c) + h

OpenStudy (plainntall):

1. Remember or Lookup the general form of a sine function 2. Substitute the numbers for the letters. If Amplitude is a and it is equal to 7 then asin(bx-c) + d is 7sinx(bx-c) + d continue substituting the numbers for the letters